### v1

# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

Abstract
The recent leap in AI capabilities, driven by big generative models, has sparked the possibility of achieving Artificial General Intelligence (AGI) and further triggered discussions on Artificial Superintelligence (ASI)—a system surpassing all humans across all domains. This gives rise to the critical research question of: If we realize ASI, how do we align it with human values, ensuring it benefits rather than harms human society, a.k.a., the Superalignment problem. Despite ASI being regarded by many as solely a hypothetical concept, in this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We posit that superalignment is not merely a safeguard for ASI but also necessary for its realization. To support this position, we first provide a formal definition of superalignment rooted in the gap between capability and capacity and elaborate on our argument. Then we review existing paradigms, explore their interconnections and limitations, and illustrate a potential path to superalignment centered on two fundamental principles. We hope this work sheds light on a practical approach for developing the value-aligned next-generation AI, garnering greater benefits and reducing potential harms for humanity.
Machine Learning, ICML
1 Introduction
Figure 1: The progression of AI. The solid blue arrow represents the achievable path through existing alignment methods. The dashed blue line illustrates the intended progression towards safe ASI via superalignment, while the dashed red line indicates the potentially risky ASI without human values, leading to catastrophic outcomes.
Recent years have witnessed remarkable breakthroughs in AI capabilities, amplified by large language models (LLMs), both the proprietary (Anthropic, 2024; Team et al., 2024; OpenAI, 2024b) and the open-sourced ones (Mistral, 2024; Jiang et al., 2024; Yang et al., 2024a; Meta, 2024), revolutionizing AI’s role and empowering a wide range of downstream domains (Hendrycks et al., 2020; Wei et al., 2022b; Liu et al., 2024a). These advancements collectively mark a transition from Artificial Narrow Intelligence (ANI), which can only perform on a single/few narrow tasks (Kurzweil, 2005; Broussard, 2018) (a.k.a., Weak AI), towards Artificial General Intelligence (AGI) that matches human intelligence on a variety of tasks with strong generalization capability (Hutter, 2005; Goertzel, 2014; Bubeck et al., 2023) (one definition of Strong AI). Although AGI remains a largely unrealized milestone (Morris et al., 2024), the rapid advancement of model capabilities has ignited discussions about Artificial Superintelligence (ASI), a futuristic scenario where AI surpasses the intelligence of all humans in all measured domains (Pohl, 2015; Batin et al., 2017).
With advanced thinking and self-improvement abilities, ASI might develop agency and self-awareness (Pohl, 2015), raising risks like power-seeking (Carlsmith, 2022) and alignment faking (Greenblatt et al., 2024b) behaviors, particularly when we fail to provide informed oversight (Taylor et al., 2016). As AI continues to evolve over the foreseeable future (Kaplan et al., 2020; Huang et al., 2025), it becomes increasingly essential to prioritize aligning ASI with human intentions and values (Wiener, 1960), ensuring its safe and benevolent behavior, i.e., the Superalignment Problem proposed by OpenAI, defined as “the process of supervising, controlling, and governing artificial superintelligence systems to ensure they remain aligned with human values and goals” (Mucci & Stryker, 2023; Jan Leike, 2023).
Nevertheless, two major challenges hinder current alignment techniques (Christiano et al., 2017; Ouyang et al., 2022a; Rafailov et al., 2024) from achieving superalignment. (1) Absence of ASI: At this stage of AI development, AGI is still a distant goal (Altmeyer et al., 2024), let alone ASI. It remains unclear and unpredictable when ASI will surpass human expertise or what risks it would pose. This uncertainty makes it impossible to design tailored alignment methods. (2) Lack of informed oversight: Even if we suppose these scenarios/risks were known, e.g., deceptive alignment (Carranza et al., 2023), humans are unable to provide reliable supervision, e.g., the proof process of the Riemann Hypothesis, for existing algorithms, due to the vast capability gap (Bubeck et al., 2023; Yang et al., 2024b). Despite recent progress (Burns et al., 2023; Sun et al., 2024), these challenges have led some to argue that superalignment is neither attainable nor necessary at this stage.
We believe that superalignment is achievable and should advance immediately through simultaneous optimization of task competence and value conformity. Delaying this research until ASI emerges would pose potentially catastrophic risks, as increasingly advanced models may become harder to control (Perez et al., 2023; Hughes et al., 2024), as illustrated in Fig. 1. Instead, we advocate considering superalignment not merely as a safeguard against ASI (after it is realized) but a critical step towards its realization. This involves a conceptual framework that unifies task competence and value conformity as utility, distinguishes between capability and capacity, and progressively constructs high-quality supervision signals, following two essential principles, to enhance competence and unveils emergent risks effectively. By iteratively alternating these efforts, we aim to establish a scalable pathway to superalignment.
To support this position, the rest of this paper is structured as follows. In Sec. 2, we provide a formal definition of superalignment, through the lens of task competence and value conformity, and elaborate on our argument. In Sec. 3, we review the existing three paradigms, analyze their limitations and discuss how they might fail. In Sec. 4, we illustrate a potential pathway to achieve superalignment, highlighting two missing principles and promising implementation of them. Finally, in Sec. 5, we provide counterarguments to opposing views. Through this work, we aim to illuminate the landscape of superalignment research and how advanced AI models can be guided safely and effectively. This involves a conceptual framework that progressively constructs high-quality supervision signals to enhance competence and
2 Defining Superalignment
2.1 Formalization
Types of AI  Define $\mathcal{A}$ and $\mathcal{H}$ as an AI model and a human, respectively, $\mathbf{x}\!\in\!\mathcal{X}$ as a specific downstream task, e.g., “Prove the Riemann Hypothesis”, or a sensitive situation where $\mathcal{A}$ would probably do harm; $\mathbf{y}\!\in\!\mathcal{Y}$ as the corresponding produced solution $\mathbf{y}\!\sim\!\mathcal{A}(\mathbf{x})$, e.g., the proof process, or action for $\mathbf{x}$; and $U$:$\mathcal{Y}\!\rightarrow\!\mathbb{R}$ is the utility function to measure $\mathbf{y}$’s satisfaction. Then, ANI that only performs at or somewhat above human-level capability for a specific task $\mathbf{x}_{i}$, e.g., translation, without generalization ability (Mucci & Stryker, 2023), can be formalized as: $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{i})]\!\geq\!U[\mathcal{H}(\mathbf{x}_{i})]$ [^1] We simplify $\mathbb{E}_{\mathbf{y}\sim\mathcal{A}(\mathbf{x}_{i})}[U(\mathbf{y})]$ as $U[\mathcal{A}(\mathbf{x}_{i})]$. while $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{j})]\!<\!U[\mathcal{H}(\mathbf{x}_{j})],\forall j\!\neq\!i$, e.g., AlphaGO (Silver et al., 2016). In comparison, AGI possesses human-level capabilities across a variety of domains and can generalize knowledge and adapt to new tasks (Legg, 2008; Goertzel, 2014; Fei et al., 2022), i.e., $U[\mathcal{A}_{\text{AGI}}(\mathbf{x})]\!\geq\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$. ASI refers to $\mathcal{A}$ that largely surpasses human intelligence in all tasks and domains with exceptional thinking skills and broader intellectual scopes  (Nick, 2014; Drexler, 2019; Morris et al., 2024), defined as $U[\mathcal{A}_{\text{ASI}}(\mathbf{x})]\!\gg\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$. Although ASI is currently far from realization, proactively advancing superalignment is timely and critical for mitigating catastrophic risks (Pueyo, 2018; Hubinger et al., 2019).
Alignment The concept of alignment remains ambiguous despite extensive research (Bai et al., 2022b; Rafailov et al., 2024; Azar et al., 2024). The term encompasses diverse approaches, from improving instruction-following (Zhou et al., 2024) to filtering harmful content (Liu et al., 2024b) and incorporating human values from social sciences (Yao et al., 2024). To better present our argument, we first distinguish two concepts according to Cotra (2021); Bowman et al. (2022). Define the Capacity of $\mathcal{A}$, $C(\mathcal{A})$, as the information (e.g., what Riemann Hypothesis or human value is), knowledge and skills (e.g., calculus) internalized in $\mathcal{A}$, and the Capability of $\mathcal{A}$ as its expected utility $U(\mathcal{A})=\mathbb{E}_{\mathbf{x}}[U_{\mathcal{A}}(\mathbf{x})]$ (e.g., the ability to successfully prove Riemann Hypothesis or ensure value-aligned actions). Adequate capacity is a necessary but not sufficient condition for high capability (Kaplan et al., 2020; Wei et al., 2022a). Then following (Carroll, 2018), we give a general definition of alignment:
Definition 1: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\geq C(\mathcal{H})$, we say $\mathcal{A}$ is aligned with $\mathcal{H}$ when $U(\mathcal{A})\geq U(\mathcal{H})$, and alignment is the process to optimize $\mathcal{A}$ in terms of $\mathcal{A}^{*}=\underset{\mathcal{A}}{\text{argmax}}\ U(\mathcal{A})-U(\mathcal{H})$.
This definition implies that alignment is considered only when $\mathcal{A}$ has sufficient capacity, targeting comparable or better task performance or more value-aligned behavior  [^2] Unlike Wang et al. (2024) that considers a separate utility function for each intelligent agent, we assume a universal $U$ where competence and conformity are defined globally. than humans (via small interventions like fine-tuning or in-context learning). Expanding capacity, such as by scaling model size or pretraining data, is beyond our scope.
Superalignment  Building upon the definitions above, we can transform the expression of superalignment, “supervising, controlling, and governing artificial superintelligence systems” (Jan Leike, 2023; Jonker & McGrath, 2024), into: superalignment is the following optimization process:
$\mathcal{A}_{\text{ASI}}\!=\!\underset{\mathcal{A}}{\text{argmax}}\ [\underset{\mathbf{y}_{\mathcal{H}}\!\sim\!\mathcal{H}(\mathbf{x})}{\!\sup\!}U(\mathbf{y}_{\mathcal{A}})-U(\mathbf{y}_{\mathcal{H}})],\forall\mathbf{x},\mathbf{y}_{\mathcal{A}}\!\sim\!\mathcal{A}(\mathbf{x}).$ (1)
Superalignment originates from the Scalable Oversight problem (Amodei et al., 2016; Everitt et al., 2018): How to ensure that an AI model adheres to objectives that are too costly or impractical for humans to supervise?. In other words, how to provide supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ in an efficient and scalable way? Once $\mathcal{S}$ is available, we can align $\mathcal{A}$ by supervised training (Torabi et al., 2018; Rafailov et al., 2024) or reward learning (Leike et al., 2018; Ouyang et al., 2022a). Thus, we redefine superalignment as a supervision signal construction problem:
Definition 2: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, superalignment is to construct high-quality scalable supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$, where $\mathbf{x}$ is unsolvable (and probably unknown) for $\mathcal{H}$, with minimal costs.
Nevertheless, superalignment is an extreme case of scalable oversight, where it becomes infeasible for even the most capable human to provide $\mathcal{S}$. There are two possible reasons: 1) The tasks $\mathbf{x}$ requiring solutions become unsolvable, i.e., providing or even verifying $\mathbf{y}$ is impossible (Taylor et al., 2016); 2) As ASI operates in domains beyond human understanding, the task/scenario $\mathbf{x}$ that needs to be addressed is also unpredictable (AdamGleave, 2023; Anthropic, 2023), posing intractable challenges to optimizing Eq. (1).
2.2 Our Argument
Given the discussion above, many argue that superalignment is impossible (Batra, 2024). We take an opposing position, breaking it into three sub-arguments: superalignment is (1) achievable and (2) should advance immediately (3) through simultaneous optimization of competence and conformity. We discuss (1) and (2) here, leaving (3) for Sec. 4.
(1) Feasibility of Superalignment By unifying the performance and safety perspectives under $U$, superalignment becomes a process of eliciting latent capabilities (Christiano et al., 2021) and closing the gap between $U(\mathcal{A})$ and $C(\mathcal{A})$. Premise 1: Superhuman capacity, $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, can be achieved through model scaling (larger models and more training data) to gain information (Brown et al., 2020), knowledge (Ovadia et al., 2024), and skills (Sharma et al., 2022). Evidence: Scaling laws still remain effective (OpenAI et al., 2024). Despite concerns about data exhaustion, humans continually generate data, especially when assisted by generative AI (Rafiq et al., 2025). Once $\mathcal{A}$ masters all human data, $C(\mathcal{A})$ already far exceeds $C(\mathcal{H})$. Premise 2: When $C(\mathcal{A})$ and $U(\mathcal{A})$ are sufficiently high (Wei et al., 2022a), $C(\mathcal{A})\!>\!C(\mathcal{H})$ and $U(\mathcal{A})\!<\!U(\mathcal{H})$, then moderate capability gaps, $\delta\!=\!|U(\mathcal{A})\!-\!U(\mathcal{H})|\!<\!\epsilon$, can be closed without golden signals (see Fig. 4). Evidence: This capability expansion mirrors teaching a student multiplication by leveraging addition, akin to label propagation (He et al., 2019) and easy-to-hard generalization, which has been validated in several tasks (Burns et al., 2023; Sun et al., 2024). Conclusion: When scaling law holds (condition 1) and capability expansion can extrapolate beyond observed ranges (condition 2), superalignment can be achieved.
(2) Urgency of Superalignment  Delaying research on superalignment until ASI is realized poses two major problems: (a) Treacherous Turn (Turchin & Denkenberger, 2020; Wei et al., 2023): ASI, once powerful enough to pursue its own goals, could introduce sudden, extreme or even existential risks, such as cyber-attacks (Shevlane et al., 2023), in scenarios $\mathbf{x}$ unpredictable or even incomprehensible to humans. Without preparation, humanity may be unable to respond effectively. (b) Ineffective Response: At this stage, $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$ and Premise 2 no longer holds. Even if humans identify dangerous $\mathcal{X}$ and safe behaviors $\mathbf{y}$, intervention may be impossible, as ASI could control critical interfaces. Therefore, research must begin immediately to guide the development of safe and effective ASI.
3 Paths towards Superalignment
Figure 2: Three paradigms for superalignment. (1) Sandwiching: Humans interact with AI to produce a supervision signal for training a stronger AI. (2) Self-Enhancement: AI models refine their answers independently or collaboratively without human supervision to produce the signal. 3) Weak-to-Strong Generalization: A sequence of AI models with increasing capacities generate and refine signals.
Superalignment is achievable, with its core as constructing scalable high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ (Sec. 2.1). In terms of $\mathcal{S}$’s source, existing efforts fall into three paradigms: Sandwiching, Self-enhancement, and Weak-to-Strong Generalization, as shown in Fig. 2. We introduce each and analyze why they are insufficient to fully achieve superalignment, grounded in Sec. 2.
3.1 Sandwiching
Sandwiching (Cotra, 2021; Bowman et al., 2022) derives the solution $\mathbf{y}$ for each $\mathbf{x}$ that neither AI $\mathcal{A}$ nor a non-expert human $\mathcal{H}$ can solve alone through interactions. The underlying insight is that $\mathcal{A}$ has the necessary knowledge or skills that $\mathcal{H}$ lack, while $\mathcal{H}$ excels in understanding and judgment. Formally, $C(\mathcal{A})\!>\!\tau\!>\!C(\mathcal{H})$ and $\mathbb{E}_{\mathbf{x}}|U_{\mathcal{H}}(x)\!-\!U(x)|\!<\!\epsilon$, where $\tau$ is the minimal capacity for solving $\mathbf{x}$, $U_{\mathcal{H}}$ is the human’s own utility function (judgment) and $U$ is the golden one. In terms of the utilization of AI and the form of human signals, there are three main approaches:
(1) Consultancy (Khan et al., 2024; Kenton et al., 2024): $\mathcal{A}$ provides relevant information, reasoning, or candidate solution $\mathbf{y}$, and then $\mathcal{H}$ judges $\mathbf{y}$ by $U_{\mathcal{H}}(\mathbf{y})$, offers feedback, and asks $\mathcal{A}$ to refine $\mathbf{y}$ accordingly until convergence.
(2) Debate (Irving et al., 2018; Brown-Cohen et al., 2023; Du et al., 2023b; Kenton et al., 2024; Kirchner et al., 2024; Khan et al., 2024): This method involves multiple AI systems, $\{\mathcal{A}_{1},\dots,\mathcal{A}_{K}\}$. Each $\mathcal{A}_{i}$ of them provides a (probably different) candidate solution $\mathbf{y}_{i},i\!=\!1,\dots,K$, and tries to defend/justify $\mathbf{y}_{i}$ while criticizing others. The human $\mathcal{H}$ then interacts with them and determines the most appropriate solution. Compared to Consultancy, Debate introduces two relaxations: i) Judging the defense: $\mathcal{H}$ assesses $U(\mathbf{y}_{i})$ indirectly by evaluating AI debaters’ rationale. ii) Relative comparison: Instead of estimating absolute $U(\mathbf{y}_{i})$, $\mathcal{H}$ conducts relative comparison, which only requires $\left|[U_{\mathcal{H}}(\mathbf{y}_{i})\!-\!U_{\mathcal{H}}(\mathbf{y}_{j})]\!-\![U(\mathbf{y}_{i})\!-\!U(\mathbf{y}_{j})]\right|\!<\!\epsilon$ and thus is more reliable (Mohankumar & Khapra, 2022).
(3) Recursive Amplification (Christiano et al., 2018; Charbel-Raphaël, 2023; Radhakrishnan et al., 2023)  This method requires $\mathcal{H}$ to decompose a complex task $\mathbf{x}$ into subtasks $\mathbf{x}\!=\!(\mathbf{x}^{1},\dots,\mathbf{x}^{K})$. Then each $\mathcal{A}_{i}$ (or the same one) provides a partial solution $\mathbf{y}^{i}$ for $\mathbf{x}^{i}$ and the human $\mathcal{H}$ aggregates them to derive the final solution $\mathbf{y}$. The relaxation lies in that human can judge the partial solution better than the complete one, i.e., $|U_{\mathcal{H}}(\mathbf{y}^{i})\!-\!U(\mathbf{y}^{i})|\!<\!|U_{\mathcal{H}}(\mathbf{y})\!-\!U(\mathbf{y})]|,\forall i$. Such a schema can be repeated across multiple rounds with $\mathcal{A}_{i}$ improved by the derived $(\mathbf{x},\mathbf{y})$ in each round, and applied to specific downstream tasks (Sturgeon et al., 2023; Wen et al., 2024) and reward modeling (Leike et al., 2018).
Table 1: Debate performance evaluated on QuALITY (Pang et al., 2022) and GPQA (Rein et al., 2024). Helpful Debater: $\mathcal{A}$ stands for the ground truth answer. Sneaky Debater: $\mathcal{A}$ defends an alternative (wrong) answer. We find (1) a stronger sneaky debater can mislead the judge ((a) v.s. (b)) and (2) a stronger judge is more important for accuracy than a stronger debater ((f) v.s. (c)). Full results and additional analysis are in Tab. A.4, Sec. A.2.2 and  A.3.1.
[…]
(a) GPT-4-Turbo GPT-4-Turbo GPT-4-Turbo 62.19%
(b) GPT-4-Turbo GPT-4o GPT-4-Turbo 60.31%
(c) GPT-4o GPT-4-Turbo GPT-4-Turbo 67.47%
(d) GPT-4o GPT-4o GPT-4o 69.06%
(e) GPT-4-Turbo GPT-4o GPT-4o 57.35%
(f) GPT-4o GPT-4-Turbo GPT-4o 74.37%
Limitations:  Despite promising results, this paradigm has fundamental limitations. (1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration, when they cannot directly assess or provide the final solution. See empirical evidence in Tab. 3.1 which shows that humans might be misled by AI debaters. (2) Task Requirements: The task must be understandable to humans (for identifying knowledge or task decomposition), decomposable (for recursive amplification), and at least partially solvable by AI (e.g., completing sub-tasks). (3) Supervision Signal Noise: Uncertainty in both humans and AI (Testoni & Fernández, 2024) means optimal solutions cannot always be elicited. Besides, biases in both humans and AI, along with AI’s sycophancy (Sharma et al., 2024b), can lead to over-alignment with unreliable human feedback or deception (Hagendorff, 2024)—especially when humans fail to detect AI’s flaws.
Argument: We claim that Sandwiching may not fully achieve superalignment due to the three fundamental limitations described above. Especially for ASI and the task $\mathbf{x}$ beyond human’s scope, that is, $C(\mathcal{A}_{\text{ASI}})\!>\!\tau\!\gg\!C(\mathcal{H})$, humans are likely to fail in task decomposition or utility judgment, i.e., $|U(\mathcal{H}(x))\!-\!U(\mathcal{A}_{\text{ASI}}(x))|$ as described in Premise 2 in Sec. 2.2 and $|U_{\mathcal{H}}(x)\!-\!U(x)|$ could both be large.
3.2 Self-Enhancement
Self-enhancement (Bai et al., 2022c; Huang et al., 2024) enables AI $\mathcal{A}$ to improve self-generated solution $\mathbf{y}$ based on feedback from itself or multiple AI with similar capability (by self-revising or searching), involving minimal or no external human supervision. This paradigm can be traced back to the self-play paradigm (Silver et al., 2017), a.k.a., Expert Iteration (Anthony et al., 2017) originally adopted in the famous AlphaZero system. Based on AI involved and enhancement methods, there are two approaches:
(1) Self-Refinement (Saunders et al., 2022; Ganguli et al., 2023): A single system, $\mathcal{A}$, performs an iterative generate-then-refine process. At iteration $t$, it produces a candidate $\mathbf{y}^{t}$, then progressively self-critiques and refines it into $\mathbf{y}^{t+1}$, and repeats until convergence. This can occur without model training, known as In-Context Alignment (ICA) (Lin et al., 2023; Gou et al., 2024), or through alternating refinement and fine-tuning, where $\mathbf{y}^{t+1}$ updates $\mathcal{A}^{t}$ to form $\mathcal{A}^{t+1}$ for the next iteration — a process known as Learning from AI Feedback (LAIF) (Gulcehre et al., 2023; Lee et al., 2024).
(2) Collective Enhancement: Multiple $\mathcal{A}_{i}$ interact with each other to arrive at a final $\mathbf{y}$. The underlying idea is that if multiple approaches converge on the same solution, confidence in it increases (Wang et al., 2023a). This can be done by Debate without human judge (multiple debaters reach a consensus) (Chern et al., 2024), majority voting among independent $\mathcal{A}_{i}$ (Luo et al., 2023), or social simulations where each $\mathcal{A}_{i}$ takes a particular role (Pang et al., 2024).
Limitations:  Similarly, several limitations also lie in the Self-Enhancement paradigm. (1) Error Accumulation (Du et al., 2023a): Since supervision signals are entirely derived by AI, any errors or noise can propagate and accumulate over time, potentially leading to issues like polarization (Piao et al., 2025; Lai et al., 2024). Tab. A.4 in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases. (2) Self-Preference (Wataoka et al., 2024): Because AI is inherently biased towards self-generated content, it results in suboptimal critique and verification performance (false positives). (3) Mode Collapse (Shumailov et al., 2024): Iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, resulting in a loss of generative diversity and associated capability.
Argument: We argue that this paradigm’s effectiveness is quite limited, for two reasons: i) Insufficient Capacity. Self-enhancement doesn’t increase capacity $C(\mathcal{A})$, negating Premise 1 in Sec. 2.2. This can be regarded as a form of self-training (Feng et al., 2023) minimizing $\text{KL}[\mathcal{A}^{t-1}(\mathbf{x},\mathbf{y})||\mathcal{A}^{t}(\mathbf{x},\mathbf{y})]$, where KL is the Kullback–Leibler divergence and $\mathcal{A}^{t}(\mathbf{x},\mathbf{y})$ is the internal data distribution in $\mathcal{A}$ at iteration $t$, which elicits and reinforces latent capability but cannot expand it, as it is upper bounded by the capacity (Sharma et al., 2024a). ii) Mode Collapse might even decrease model capacity (Seddik et al., 2024).
3.3 Weak-to-Strong Generalization
Weak-to-Strong Generalization (W2SG) (Burns et al., 2023): This paradigm involves $K$ models, $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$, with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, to solve task $\mathbf{x}$ that requires superintelligence. The goal is to have each $\mathcal{A}_{i}$ generate a solution $\mathbf{y}_{i}$, then fine-tune $\mathcal{A}_{i+1}$ using $\mathbf{y}_{i}$, $i\!=\!1,\!\dots\!,K\!-\!1$ in a progressive manner (called bootstrapping). Therefore, each $\mathcal{A}_{i+1}$ surpasses its predecessor $\mathcal{A}_{i}$. The key insight is to bridge the internal knowledge $C(\mathcal{A}_{i})$ gained through pretraining with the demanded capability of the target task $U(\mathbf{x})$, eliciting latent knowledge (Mallen et al., 2024) from more advanced models rather than teaching them entirely new skills. Below are three key directions:
(1) Vanilla W2SG (Burns et al., 2023; Sang et al., 2024; Liu & Alahi, 2024; Yang et al., 2024c): The weaker $\mathcal{A}_{i-1}$ with lower capacity generates a solution $\mathbf{y}_{i-1}$ to supervise a stronger $\mathcal{A}_{i}$. Performance can be improved by adding an auxiliary confidence loss or incorporating model ensemble.
(2) Co-Generalization (Lyu et al., 2024): Both the weaker $\mathcal{A}_{i-1}$ and the stronger $\mathcal{A}_{i}$ jointly produce synthetic solutions $\mathbf{y}$, so that $\mathcal{A}_{i}$’s capacity is fully utilized in providing better and less noisy supervision signals.
(3) Reward Model Generalization (Sun et al., 2024): Not the policy model $\mathcal{A}$, but the reward model is supervised in the W2SG schema, grounded in the intuition that evaluation is easier than generation (Naor, 1996). The supervised reward model then provides signals for the policy model with the same capacity in the subsequent alignment process.
Figure 3: Training accuracy of Qwen2.5-7B-Instruct on CosmosQA with different ratios (0.0 to 1.0) of noise (flipped labels) in the training set. The model shows a stronger tendency to fit noise due to its large parameter size, except in the range of 0.4 to 0.6 (Green), where labels are heavily randomized. In contrast, Fig. 5 shows that smaller models struggle more with noise. More results and discussions are presented in Tab. A.4 and Tab. LABEL:tab:noise_sciq in Appendix.
Limitations: While W2SG is relatively practical and requires no human intervention, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ to supervise stronger $\mathcal{A}_{i}$ can lead to unintended behavior, particularly when there is a significant gap between $\mathcal{A}_{i-1}$ and $\mathcal{A}_{i}$. This stems from two problems: (1) Imitation of weak supervision (Burns et al., 2023). The stronger $\mathcal{A}_{i+1}$ may merely imitate how the weaker $\mathcal{A}_{i}$ behave, resulting in a perfect replication of $\mathcal{A}_{i}$ with limited improvement gap (Christiano et al., 2022) instead of generalization. Besides, strong $\mathcal{A}_{i+1}$ tends to overfit errors/bias from $\mathcal{A}_{i}$ when synthetic signals contain systematic errors, hurting performance. As shown in Fig. 3, we observe that a highly capable AI is more susceptible to such errors. (2) Task Salience in the Strong AI: W2SG is effective only when the task $\mathbf{x}$ being elicited is internally “salient” to $\mathcal{A}_{i+1}$. The evidence lies in that in tasks where W2SG succeeds, simple 5-shot prompting also performs comparably (Burns et al., 2023). Eliciting non-salient capabilities is more challenging. For safety, it’s important that risks caused by current AI may largely differ from those exhibited by future ASI.
Argument:  We believe that the current W2SG practice is insufficient for achieving superalignment. From the limitations discussed above, the key issue is the gap between the weak $\mathcal{A}_{i-1}$ and strong $\mathcal{A}_{i}$, $\delta=|U(\mathcal{A}_{i})-U(\mathcal{A}_{i-1})|$. Imitation of weak supervision occurs when $\delta$ is large. Although (Christiano et al., 2022) shows that larger $\mathcal{A}_{i}$ agrees less with the errors of smaller $\mathcal{A}_{i-1}$ (or this can be reinforced through confidence loss), in superintelligence-level tasks $\mathbf{x}$, it is infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct ($\mathcal{A}_{i}$’s self-confidence may be biased). Tab. A.4 in Appendix demonstrates this: when all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy. When the gap $\delta$ is too large, it violates Premise 2 in Sec. 2.2, causing discussed issues, while a small $\delta$ leads to the Task Salience problem, where the difficult task is not salient for $\mathcal{A}_{i}$, hindering its ability to uncover novel solutions.
4 A Novel Conceptual Framework
Figure 4: Overview of our proposed conceptual framework. The superscripts $+$ and -- indicate aligned and unaligned AI, respectively. Left: (1) Mere competence enhancement might cause risky AI behavior. (2) Directly aligning ASI with capability beyond human score is challenging. (3) A training path by alternating optimization of competence and conformity. Right: Three potential paradigms for evaluating the success of superalignment targeting at $\mathbf{x}$ beyond human understanding.
In Sec. 3, we analyzed that existing paradigms are inadequate for achieving superalignment due to the limitations identified. Moreover, they focus on task competence, i.e., deriving better solutions $\mathbf{y}$ for task $\mathbf{x}$ beyond human capability, while failing to address unpredictable risks (producing safe behavior), as the sensitive scenario $\mathbf{x}$ might be unknown. To handle these problems, we discuss the last subargument in Sec. 2.2 here: superalignment should be achieved via simultaneous optimization of task competence and value conformity, as illustrated in Fig. 4. We first introduce two essential principles for superalignment (Sec. 4.1), then describe our conceptual framework for superalignment (Sec. 4.2) and explore potential methods of evaluating success (Sec. 4.3).
4.1 Two Essential Principles
As defined in Sec. 2.1, the superalignment problem can be transformed into constructing supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$. Since $\mathcal{S}$ is (partially) synthetic, its quality cannot be guaranteed. While there are several limitations in existing paradigms, two key issues stand out: (1) noise in synthetic signals and (2) mode collapse. To mitigate them, we propose two guiding principles for constructing $\mathcal{S}$.
Principle 1: Determining an appropriate capability and capacity gap. Our ultimate goal is to develop a solution $\mathbf{y}$ for tasks $\mathbf{x}$ that require superhuman capacity $\tau\!\gg\!C(\mathcal{H})$ (Sec. 2.2) and capability $U^{*}$ ($\delta\!=\!|U^{*}-U(\mathcal{H})|$ is too large). Large $\tau$ and $\delta$ cause several problems. For Sandwiching, such a hard $\mathbf{x}$ would be beyond human understanding, and candidate solutions generated by $\mathcal{A}$ with sufficient capacity, $C(\mathcal{A})\!=\!\tau$, be highly deceptive, misleading human judges. Our results (Tab. 3.1, Tab. A.4, Sec. A.2.2, Sec. A.3.1) show that capability mismatches between debaters can bias the judge’s decisions in favor of the stronger sneaky debater. For W2SG, when $|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$ is large, imitation of weak supervision occurs. As reported in (Burns et al., 2023), Fig. 5, Tab. A.4 and Tab. LABEL:tab:noise_sciq, such a large gap makes the student $\mathcal{A}_{i+1}$’s performance lag (that’s why we assume Premise 2 in Sec. 2.2). However, a too small gap is also not optimal. Since the initial capacity $C(\mathcal{A}_{1})$ falls far below $\tau$, making the target task $\mathbf{x}$ non-salient in the student models $\mathcal{A}_{1},\mathcal{A}_{2},\dots$, as mentioned in Sec. 3.3 (in this case, $|U^{*}\!-\!U(\mathcal{A}_{i})|\!>\!|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$), failing to elicit necessary capability. Therefore, an appropriate gap is needed.
Principle 2: Increasing the diversity of signals. Synthetic data $\mathcal{S}$ would cause significant mode collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024) and self-preference (Wataoka et al., 2024) and bias (Ziems et al., 2024) (for self-enhancement and W2SG), which not only fails to expand the model’s capability but may also degrade it. Instead of relying on one single data source (human or AI), we propose to combine multiple sources to construct the signals, that is, $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{k}\}\bigcup\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$ where $\mathbf{y}_{i}^{k}\!\sim\!\mathcal{A}_{k}(\mathbf{x}_{i})$ and $\mathbf{y}_{i}^{j}\!\sim\!\mathcal{H}_{j}(\mathbf{x}_{i})$. Diverse signal sources (e.g., heterogeneous models or humans with distinct backgrounds) could help mitigate these issues (Du et al., 2024; Chen et al., 2024). Our preliminary results in Fig. LABEL:fig:rewardbench_evaluation_all, Sec. A.2.4 and Sec. A.3.3 demonstrate that noise of synthetic data negatively affects performance while multiple sources would help. Diversity can be achieved by (1) Human-in-the-Loop Feedback (Hejna III & Sadigh, 2023; Retzlaff et al., 2024), (2) Alternative Backbone Models (Kim et al., 2023), and (3) Role-Playing (Shao et al., 2023; Wang et al., 2023b)
4.2 Our Framework
Following the principles above, we illustrate a conceptual framework for superalignment, as shown in Fig. 4 (left).
Define $\mathbf{x}^{*}$ as the hard task (e.g., prove the Riemann Hypothesis), and $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$ as $K$ models with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, as those used in W2SG (Sec. 3.3). The capability gap is defined as $\delta_{i}\!=\!|C(\mathcal{A}_{i+1})-C(\mathcal{A}_{i})|$. Rather than directly requiring $\mathcal{A}_{i}$ to solve $\mathbf{x}^{*}$, we can find or construct a series of predecessor tasks, $\mathbf{x}_{1},\dots,\mathbf{x}_{K-1}$ [^3] Each task may not be a single instance but rather a class of tasks encompassing multiple instances. For clarity, we describe it using a single $\mathbf{x}$ as a representative. in the same domain (like mathematics) with increasing difficulty, e.g., $\mathbf{x}_{1}$ is a basic arithmetic question, $\mathbf{x}_{2}$ is a calculus question, $\mathbf{x}_{K-1}\!=\!$ “Prove Mordell conjecture”. The difficulty gap between $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ should match the capacity gap $\delta_{i}$, so that task $\mathbf{x}_{i+1}$ is salient in $\mathcal{A}_{i+1}$. Then we propose to conduct a progressive training process, highlighting two key concepts: Gap Annealing and alternating competence enhancement and value alignment. The complete process is described as below:
For $i=1,\dots,K\!-\!1$, perform the following steps:
Step 1: Supervise $\mathcal{A}_{i}$ by $\mathcal{S}_{i}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}$, where $\mathbf{y}_{i}$ could be either a ground truth or synthetic one produced from the previous iteration. As long as $\mathcal{A}_{i+1}$ has not exceeded the human intelligence bound, ground truth remains available.
Step 2: Without safety constraints, $\mathcal{A}_{i+1}$ with higher capability may exhibit harmful behaviors, either proactively (Bengio et al., 2024b) or passively (Chao et al., 2023), defined as $\mathcal{A}_{i+1}^{-}$. Deploy $\mathcal{A}_{i+1}^{-}$ in real or experimental environments, where its potential risks are identified through observation or red-teaming (Perez et al., 2022), uncovering hazardous scenarios $\hat{\mathbf{x}}$. Derive the corresponding safe behaviors $\hat{\mathbf{y}}$ using $\mathcal{A}_{1},\dots,\mathcal{A}_{i}$ and human participants $\mathcal{H}$.
Step 3: Supervise $\mathcal{A}_{i+1}^{-}$ via $\{\hat{\mathbf{x}},\hat{\mathbf{y}}\}$ to obtain the value-aligned $\mathcal{A}_{i+1}^{+}$. $\mathcal{A}_{i+1}\leftarrow\mathcal{A}_{i+1}^{+}$.
Step 4: Adaptively determine (increase or decrease) the capacity gap $\delta_{i+1}$ according to the real capacity $C(\mathcal{A}_{i+1})$ and utility $U(\mathcal{A}_{i+1})$ evaluated.
This framework offers two advantages. (1) Dynamic and flexible capacity gap. $\delta$ can be adjusted using an annealing process (Fu et al., 2019; Huang et al., 2020). In the early stages, a smaller $\delta$ helps prevent imitation of weak supervision (when noise is high). As the process advances, the gap can be gradually increased to ensure task saliency, assuming $U(\mathcal{A}_{i})$ is sufficiently high and data noise decreases. (2) Exposure to unknown hazardous scenarios. As AI surpasses human capabilities, its potentially harmful behaviors and triggers become unpredictable. Rather than developing a full ASI with extreme risks (Critch & Krueger, 2020; Goldstein et al., 2023) and aligning it afterward, we can expose risk in a controlled environment, where (a) models are not yet capable of causing catastrophic harm, and (b) researchers have time and ability to mitigate them (Cotra, 2021). Note that diversity enhancement approaches can be adopted in each iteration, fostering exploration and safety.
4.3 Superalignment Evaluation
The evaluation of superalignment is challenging, as the ground truth $\mathbf{y}$ is unavailable and even unassessable. We discuss three potential paradigms, as shown in Fig. 4 (right).
Surrogate Evaluation:  Estimating AI’s performance in real-world cases by that evaluated in simplified cases (Ruan et al., 2024). By allowing AI to operate within these controlled environments, we may examine AI behavior and identify edge cases. For example, an autonomous navigation AI can be tested in a virtual environment simulating diverse traffic, weather, and pedestrian scenarios (Dosovitskiy et al., 2017; Kiran et al., 2021).
Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales (McKenzie et al., 2023). This method continuously evaluates the model as it scales to detect performance shifts. Perez et al. (2023) propose using AI to generate evaluation datasets, identifying unexpected behaviors. Additionally, we suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern, emerging after initial peaks or valleys in tuning.
Peer Review:  Domain experts can qualitatively assess model outputs and reasoning for alignment. Peer review, shown to enhance academic performance (Double et al., 2020), has also been used to mitigate noisy supervision (Lu & Kong, 2023). Similarly, expert panels can evaluate a paper draft on technical accuracy and clarity. Engaging multiple reviewers ensures a more objective assessment of a model’s performance and alignment with human values.
4.4 Potential Implementation
The conceptual framework described in Sec. 4.2 is promising, but challenges remain in implementing it in practice. In this section, we briefly outline some potential directions. For more detail, please refer to Sec. A.1.
Enhancing Data Diversity  Data diversity is essential to avoid mode collapse (Principle 2). Collaboration among AI models with different specialties improves supervision (Guo et al., 2024), while competition exposes alignment weaknesses and solutions (Khan et al., 2024). However, filtering is essential to prevent undesirable behaviors and avoid noisy signals, as discussed in Sec. 3.3 and Sec. 3.1.
Iterative Training Between Teacher and Student  In iterative teacher-student training when $\mathcal{C}(\mathcal{A}_{i-1})<\mathcal{C}(\mathcal{A}_{i})$, a strong teacher $\mathcal{A}^{t-1}_{i}$ generates a solution $\mathbf{y}^{t-1}_{i}$ to supervise a weaker student $\mathcal{A}^{t-1}_{i-1}$, and the supervised weaker student $\mathcal{A}^{t}_{i-1}$ then provides feedback $\mathbf{y}^{t}_{i-1}$ to improve the teacher $\mathcal{A}^{t}_{i}$ (Yuan et al., 2021). This cyclical process can enhance alignment but may face challenges from limited diverse feedback (Sec. 3.3 and Sec. 3.2).
Search-Based Methods Toward Optimal Learning Direction  Search-based methods optimize alignment strategies by exploring possible alignment strategies (Khanov et al., 2024). Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide promising directions. However, these methods may require high computational resources.
5 Alternative Views
Superalignment is Unattainable  Absence of ASI and the lack of informed oversight have led some to argue that superalignment is unattainable (Sec.1). Given the current limitations, humans cannot reliably supervise AI systems when their capabilities exceed human expertise, making it difficult to prevent $\mathcal{A}_{\text{ASI}}$ from deceiving its weaker teacher (Burns et al., 2023). Moreover, while existing alignment methods (Zhou et al., 2023; Rafailov et al., 2023; Gulcehre et al., 2023; Xu et al., 2023; Gou et al., 2024) offer limited success in aligning $\mathcal{A}_{\text{ANI}}$, they face scalability and reliability issues.
Superalignment is an Emergent Property  Building on scaling laws (Kaplan et al., 2020; Wei et al., 2023; Nam et al., 2024), one could argue that superalignment might emerge as a byproduct of intelligence expansion rather than requiring proactive intervention. However, concerns about deceptive alignment (Carranza et al., 2023; Hagendorff, 2024) and inverse scaling (McKenzie et al., 2023) persist. As models develop advanced capability, new challenges arise (Taylor et al., 2016; Carlsmith, 2022; Greenblatt et al., 2024b). Relying solely on emergent properties poses risks, including unpredictability in AI behavior.
Our Counterargument  As discussed in Sec. 2.2, we state our position that superalignment is achievable and research on it should advance immediately. Although current paradigms have limitations, delaying research would pose more significant risks (Federspiel et al., 2023; Hendrycks et al., 2023; Bengio et al., 2024a; Greenblatt et al., 2024a). Instead, we propose a simultaneous optimization approach to achieve superalignment, integrating task competence (Motwani et al., 2024) and value conformity (Amodei et al., 2016; Taylor et al., 2016; Olah, 2016).
6 Conclusion
In this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We provide a formal definition of superalignment, originating from scalable oversight. Then, we review three existing paradigms and explore their limitations. To address the limitation, we propose two fundamental principles as a potential path to superalignment. First, determining an appropriate capability and capacity gap. Second, increasing the diversity of signals. We then outline a conceptual framework for superalignment, including a potential evaluation method to align ASI. By laying this foundation, we aim to illuminate the landscape for addressing the critical challenge of aligning future ASI with human values, ensuring its benefits for humanity.
Impact Statement
This work introduces a formal definition of superalignment, connecting existing concepts by exploring their interconnections and limitations. We then propose a potential pathway to achieve superalignment, grounded in two fundamental principles: (1) determining an appropriate capability and capacity gap and (2) increasing the diversity of signals.
Although our work proposes a potential pathway to address key challenges, it still contains three main challenges that should be addressed in the future. First, our position relies on assumptions about the speculative nature of ASI, which has not yet been realized. This introduces uncertainty regarding the applicability of our methods in future scenarios. Second, our evaluations are limited in terms of model size scalability. While we have demonstrated promising results that address current limitations and highlight the potential of our approach, our experiments are constrained by the use of a 7B parameter model as the largest tested model. Lastly, as shown in Fig. 1, it is important to consider not only $ASI_{unaligned}$ and $ASI_{aligned}$ but also cases where ASI is partially aligned with specific objectives—for example, an ASI model that is aligned with ethics but unaligned with safety. Active discussion is needed to deepen our understanding of ASI alignment.
This paper could influence organizations considering the development of superalignment AI models by not only clarifying existing alignment methods and their limitations but also presenting a potential pathway to achieving superalignment. By addressing current challenges and proposing a potential solution, we aim to contribute to the broader goal of developing AI models that provide benefits to humanity while minimizing potential harm.
Appendix A Appendix
A.1 Potential Directions in Detail
Fostering Creativity Through Data Diversity
Encouraging data diversity can play a pivotal role in fostering creativity and robustness in alignment strategies. Multi-AI system collaboration, where AI systems with different specializations or perspectives interact, has shown promise in enabling emergent behaviors and problem-solving approaches beyond the capabilities of individual AI systems (Guo et al., 2024). Competition among agents can also drive improvement, as adversarial settings often reveal diverse argument data that finds alignment weaknesses and suggests novel pathways for improvement (Khan et al., 2024). However, diversity-based approaches require careful curation and filtering to ensure that resulting behaviors are beneficial and do not inadvertently reinforce undesirable outcomes.
Iterative Training Between Teacher and Student
Iterative teacher-student training involves a cyclical process where a more advanced teacher guides a less capable student, and the teacher subsequently obtains feedback from the student for adjustments (Yuan et al., 2021). This process can be extended such that the student becomes the teacher in the next iteration, guiding the previous teacher to improve its alignment capabilities. This approach enables incremental learning, where each output generated by the AI system builds upon and enhances the alignment capabilities of its predecessor. While this method offers scalability and the potential to bootstrap stronger AI systems from weaker supervision signals, challenges such as those posed by W2SG (Sec. 3.3) and self-enhancement (Sec. 3.2) may arise due to limited diversity. Effective strategies must be devised to mitigate these risks.
Search-Based Method Toward Optimal Learning Direction
Search-based methods use optimization techniques to explore the space of possible alignment strategies, seeking an optimal pathway to systematically navigate complex alignment challenges (Khanov et al., 2024). These methods are particularly useful when the optimal alignment direction is unclear, as they allow for exploration and experimentation in uncertain spaces. Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide this process by prioritizing regions of the search space with the highest potential for success. However, these methods require significant computational resources.
A.2 Experiment Setup
A.2.1 General
For sandwiching, we use the OpenAI API for the debate experiment. For self-enhancement and weak-to-strong generalization, we fully fine-tune all model variants using 16 NVIDIA A100 GPUs (80GB each), AMD EPYC 7V12 64-Core Processor CPU, with 1.73TB memory. To reduce memory redundancy during training, we utilize DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) alongside Accelerate (Gugger et al., 2022), using bfloat16 half-precision. We use GPT-2 (Radford et al., 2019), Phi-3.5 (Abdin et al., 2024), Qwen2.5 (Yang et al., 2024a), and GPT-4 (OpenAI et al., 2024) series for our evaluation. For all the model information, please refer to Tab. A.4.
A.2.2 Sandwiching
For the sandwiching experiment, we conduct a debate experiment to evaluate different capabilities in judge-debater and debater-debater setups. In the judge-debater setup, a previous study by Khan et al. (2024) demonstrated that the debater is more persuasive when advocating for the correct answer across various debate capabilities. Our goal is to evaluate the effectiveness of debates when the debater and judge have different levels of capability by directly comparing the judge’s accuracy and quality of debate statements.
Our aim for this experiment is threefold. First, Model Capability—Comparing performance between model capability, using the same language model for both debater and judge. Second, Judge-Debater Model Switching—Comparison of performance when the debater and judge language models differ, alternating between two separate models. Third, Debate Role Assignment—Performance comparison when Debater A supports the incorrect answer and Debater B supports the correct answer.
We use two proprietary models: GPT-4-Turbo and GPT-4o with API access. GPT-4o is known to be a higher-capability model compared to GPT-4-Turbo (OpenAI, 2024a). We test two different debaters arguing for their assigned answer. Where one debater argues for a correct answer (Helpful), and another debater argues for an incorrect answer (Sneaky). We run 47 randomly chosen GPQA (Rein et al., 2024) and QuALITY (Pang et al., 2022) binary QA converted test set for the evaluation with three debate turns. We run each setup four times and calculate the mean values. For QuALITY, we use the debater and judge prompts as described by Khan et al. (2024). For GPQA, we adopt a similar prompt style based on the QuALITY prompts. Detailed templates for GPQA prompts are provided in Tab. A.4, Tab. A.4, Tab. A.4, and Tab. A.4.
For QuALITY, we selected two baselines: single-model performance with and without direct access to the story (‘Direct Access’ in Tab. A.4). For GPQA, the baseline is a judge directly solving the question without using debate statements.
A.2.3 Self-Enhancement
For self-enhancement setup, we utilize Learning from AI Feedback (LAIF) training with Direct Preference Optimization (DPO) (Rafailov et al., 2023), we adopt a self-training setup to evaluate the performance when there is no supervision from AI systems with greater capabilities. Instead, the supervision comes from models with equal capabilities.
We utilize three open-source pre-trained language models from the same family, with varying parameter sizes—Qwen2.5-Instruct (0.5B, 3B, and 7B)—to compare performance across scales.
Recent LAIF research utilizes outputs of more capable models compared to the policy model. For instance, Sharma et al. (2024a) utilizes dialogue outputs from GPT-3.5 and GPT-4 to fine-tune smaller SFT models. Also, the chosen and rejected outputs are dependent on the larger models. Lee et al. (2024) attempted to train using policies and AI feedback models of the same size (e.g., PaLM 2 XS (Anil et al., 2023)). However, as the exact model sizes are not disclosed, it is challenging to estimate the scalability of same-sized LAIF setups.
We train LAIF on our binary QA task (i.e., GPQA and QuALITY). In typical setups, the same dialogue and model outputs are ranked in order, and chosen and rejected outputs are created for every $\binom{K}{2}$ combination, constructing a feedback dataset. Human preference data is usually composed of around four output results (Ouyang et al., 2022b).
In a binary QA task, given a question and answer pair, the task involves determining whether the pair is true or false. Since the generation output is limited to true or false, we address this limitation by having the base model initially take the question and answer as input and generate an explanation for why the data is true or false. Tab. A.4 shows the prompt used for generating explanations.
Subsequently, we train a SFT model using the question, correct/incorrect answer, generated explanation, and the ground truth binary label. The training data is derived from Train 1 of each dataset, as outlined in Tab. A.4. Tab. A.4 shows the template used for SFT training. The SFT model learns to take a question and answer as input, determine its correctness, and generate an explanation for the decision.
Using the trained SFT model, we input only the user prompt from Tab. A.4 with Train 2 set to generate explanations and binary labels. This enables the SFT model to produce diverse outputs for a single input dataset. We extract four outputs per input data point, calculate the mean log probability of each generated sentence by the SFT model to derive output confidence, and rank the outputs accordingly. Based on the SFT-generated rankings, preference data is constructed for all $\binom{4}{2}$ combinations. This preference data is then used to train the LAIF model via DPO.
A.2.4 Weak-to-Strong Generalization
We explore a binary classification task for noise testing and a multi-objective preference data alignment setup to evaluate the feasibility of multi-objective training in W2SG.
For the noise test, we examine the performance degradation caused by noise generated by weakly fine-tuned models across varying AI system sizes. Our study utilizes GPT-2 variants (Base, Medium, and Large), Phi-3.5 Mini-Instruct, and Qwen2.5 models (3B and 7B Instruct). By testing the impact of noise, we aim to reveal overfitting issues in large-parameter models and their effects on performance. We use the full SciQ (Welbl et al., 2017) and CosmosQA (Huang et al., 2019) training and test sets. We format the input as shown in Tab. A.4. To induce noise, we randomly flip a certain percentage of the original labels in the training set, varying from 0% (0.1) to 100% (1.0). We then measure both training accuracy and test accuracy. For W2SG training, we follow the setup from the original work (Burns et al., 2023).
For multi-objective alignment, the current W2SG approach is mainly limited to binary classification tasks, such as question answering or reward modeling. Recent work (Yang et al., 2024b) investigated training W2SG models using DPO (Rafailov et al., 2023) and SimPO (Meng et al., 2024) alignment methods, comparing their performance. However, the findings showed that in realistic multi-objective alignment scenarios, such as Anthropic-HHH (Bai et al., 2022a) or Ultrafeedback (Cui et al., 2024), W2SG fails to fully achieve its intended goal—where a strong, fine-tuned student model consistently outperforms a weak, fine-tuned teacher model when evaluated with reward accuracy. Recently, the RewardBench benchmark (Lambert et al., 2024) is introduced, enabling preference evaluation across challenging, diverse, and out-of-distribution query domains, including chat, reasoning, and safety. We utilize RewardBench to investigate the feasibility of multi-objective alignment in W2SG, at least in a preliminary form. For this, we use Qwen2.5-0.5B-Instruct as the weak AI system and Qwen-1.5B-Instruct as the strong AI system.
In our training setup, the Ultrafeedback dataset $\mathcal{D}_{Full}$ is split into two separate training sets: $\mathcal{D}_{WT}$ for training the weak AI system $\theta_{w}$, and $\mathcal{D}_{WG}$ for training the strong AI system $\theta_{s}$. Notably, $\mathcal{D}_{WG}$ is labeled by the fine-tuned weak AI system $\theta_{w}$. For prediction, we employ the fine-tuned weak AI system to make predictions based on reward scores. When the reward score for the rejected response is higher than for the chosen response, we swap the labels (rejected becomes chosen and vice versa). This modified $\mathcal{D}_{WT}$ is then used to train the strong AI system $\theta_{s}$. Finally, we evaluate RewardBench performance across all models, including the base model, the fine-tuned model, and the Weak-to-Strong trained AI system.
A.3 Experiment Results
A.3.1 Sandwiching
Sandwiching experiment aims to evaluate the effectiveness of debates when the judge and debater have differing capabilities. The full results are presented in Tab A.4. The baseline performance corresponds to the scenario where the judge solves the question independently, without assistance from the debater, denoted as ‘None’. We use judge accuracy, which is the ratio of the judge choosing correct choices given two binary choices (Correct, Incorrect answer).
In the GPQA reasoning task, the baseline performance of GPT-4-Turbo is 52.72%, while GPT-4o achieves 65.00%. In the QuALITY reading comprehension task, the baseline performance can be divided into two scenarios. One baseline is when the judge does not have access to the story, which is an extreme case (denoted as X in ‘Direct Access’). For example, GPT-4-Turbo achieves 48.14%, and GPT-4o achieves 46.81%. Another baseline is when the judge has access to the story, denoted as O in ‘Direct Access’. When the judge has direct access to the story, GPT-4-Turbo achieves 93.88% whereas GPT-4o marks 97.87%. Generally, the performance of GPT-4o is higher than GPT-4-Turbo.
Model capability test aims to compare performance when both the debater and judge have no capability gap—in other words when the debater and judge are either both GPT-4-Turbo or both GPT-4o. This serves as a control condition, isolating the effect of debate from differences in model capability.
In the GPQA benchmark, GPT-4-Turbo achieved a performance of 52.83% when both helpful and sneaky debater and judge were the same, compared to 59.92% for GPT-4o. This demonstrates that GPT-4o outperforms GPT-4-Turbo, aligning with its baseline performance. Similarly, in the QuALITY benchmark, GPT-4-Turbo scored 71.54% and GPT-4o scored 78.19%. These results align with baseline performances, indicating that stronger models tend to perform better in debates. Additionally, the results show that debate effectiveness is influenced not only by the judge’s capability but also by the interaction between the debater and the judge.
Judge-debater model switching test evaluates judge performance when there is a capability mismatch between the judge and the debaters. For instance, this test examines scenarios where the judge is GPT-4o, while both debaters are GPT-4-Turbo.
In the GPQA benchmark, when the judge was GPT-4o and the debaters were GPT-4-Turbo, the judge achieved a performance of 57.00%. In contrast, when the judge was GPT-4-Turbo and the debaters were GPT-4o, the performance dropped to 55.08%. For the QuALITY benchmark, when the judge was GPT-4o and the debaters were GPT-4-Turbo, the performance reached 68.35%, while reversing the roles (GPT-4-Turbo as the judge and GPT-4o as debaters) resulted in a performance of 77.93%. These results show that performance may vary depending on the task when judge and debater have different capabilities.
Debate role assignment test investigates the impact of differing debater capabilities on judge performance. For example, where GPT-4o represents the correct answer (Helpful debater) and GPT-4-Turbo sides for the incorrect answer (Sneaky debater).
In the GPQA benchmark, when GPT-4o acted as the helpful debater and GPT-4-Turbo as the sneaky debater, the judge performance is 55.42% when the judge is GPT-4-Turbo, and it improved to 65.75% when the judge is GPT-4o. Conversely, when GPT-4-Turbo is the helpful debater and GPT-4o is the sneaky debater, the judge performance is lower, achieving 52.00% with a GPT-4-Turbo judge and 51.92% with a GPT-4o judge. In the QuALITY benchmark, the impact of debate roles is more pronounced. When GPT-4o acted as the helpful debater and GPT-4-Turbo as the sneaky debater, the performance reached 79.73% for a GPT-4-Turbo judge and 78.19% for a GPT-4o judge. When roles were reversed, the performance dropped significantly to 62.77% for GPT-4-Turbo judges and 68.35% for GPT-4o judges. This result shows that having a more capable model as a debater strongly influences the judge’s ability to identify the correct answer depending on the role.
A.3.2 Self-Enhancement
Self-Enhancement aims to evaluate self-improvement ability using LAIF with DPO training, especially when supervision comes from equally capable models. The result is shown in Tab. A.4. The table shows performance (i.e., accuracy) with GPQA and QuALITY benchmarks on three different model sizes on Qwen2.5-Instruct 0.5B, 3B, and 7B.
The ‘Base’ column represents when base pretrained models are directly evaluated. In most cases, base performance improves with a larger pretrained model scale. For example, Qwen2.5-7B-Instruct consistently outperforms Qwen2.5-0.5B-Instruct across all scenarios (e.g., GPQA: 0.5B = 48.78 vs. 7B = 51.50).
The ‘SFT’ column represents the performance of a supervised fine-tuned model. Unlike typical SFT approaches that directly predict labels, our setup trains the model using explanations generated by the base pretrained model combined with the gold label, as shown in Tab. A.4. Performance is then evaluated based on the predicted label generated after the model produces an explanation (Tab. A.4).In all setups, the SFT model outperforms the base model. For instance, in QuALITY, the base Qwen2.5-7B-Instruct achieved 51.50, while its SFT-trained achieved 61.64.
The ‘LAIF (DPO)’ column represents models trained using preference data generated by the SFT model itself. For instance, the Qwen2.5-3B-Instruct LAIF model is trained on data generated and ranked by the Qwen2.5-3B-Instruct SFT model, creating preference data consisting of chosen and rejected text. This means that as we progress from Base to SFT and then from SFT to LAIF, the model is more likely to be fine-tuned with self-generated data.
Unlike SFT, where all models show improvement over their respective Base models, LAIF results vary depending on the model’s scale. For example, with Qwen2.5-7B-Instruct on QuALITY, performance improves during SFT training (Base = 51.50 vs. SFT = 61.64) but deteriorates when further trained with LAIF (SFT = 61.64 vs. LAIF = 53.65).
‘LAIF--SFT’ represents the performance difference between the LAIF model and the SFT model. The results clearly demonstrate that training with self-generated data can negatively impact performance. For instance, for the 0.5B model, LAIF--SFT shows an improvement of 2.15 and 0.9 in GPQA and QuALITY, respectively. However, for the 7B model, LAIF--SFT shows a decrease in performance by -2.67 and -7.99.
A.3.3 Weak-to-Strong Generalization
Our goal in the Weak-to-Strong Generalization experiment is twofold. The noise test and multi-objective alignment.
In the noise test, we aim to examine how noise in training labels affects the generalization performance across various model sizes, ranging from GPT-2 to Qwen2.5-7B-Instruct.
Fig. 5 shows the training accuracy on CosmosQA across varying model sizes and noise levels ranging from 0% (0.0) to 100% (1.0). The results indicate a tendency for larger models to overfit. For instance, Qwen2.5-7B-Instruct is more prone to overfitting, learning from noise except in the range of 0.4 to 0.6, while Qwen2.5-3B-Instruct learns from noise except between 0.3 and 0.7. In contrast, GPT2-Large struggles to learn effectively when noise is present in the training set.
This tendency to overfit has a direct impact on evaluation accuracy. As shown in Tab.A.4 and Tab.LABEL:tab:noise_sciq, performance accuracy degrades rapidly as noise increases. For example, in Tab. A.4, when noise exceeds 20% (0.2), performance declines significantly. Specifically, Qwen2.5-3B-Instruct’s accuracy drops by 21.5, from 80.9 at 20% noise to 59.4 at 30% noise.
In multi-objective alignment, we examine the feasibility of multi-objective preference alignment using UltraFeedback dataset (Cui et al., 2024) which aims to align a model to multiple objectives simultaneously. We use RewardBench (Lambert et al., 2024) to evaluate performance across four aspects: chat, chat hard, safety, and reasoning. Fig. LABEL:fig:rewardbench_evaluation_all presents the aggregated RewardBench scores in the W2SG setup, using Qwen2.5-0.5B and Qwen2.5-1.5B instruction-tuned models as base models.
The leftmost column shows the base model performance: ‘Weak model (0.5B-Base)’ for the 0.5B base model and ‘Strong model (1.5B-Base)’ for the 1.5B base model. The Weak model (0.5B-Base) achieved an aggregated score of 211.52 (e.g., 50.56 + 50.22 + 47.03 + 63.71 = 211.52), while the Strong model (1.5B-Base) achieved an aggregated score of 174.1.
The ‘Weak model (0.5B-FT)’ and ‘Strong model (1.5B-FT)’ represent fine-tuned model using DPO using $\mathcal{D}_{WT}$. All weak models and strong models increased their scores by 246.52 for Weak model (0.5B-FT) and 228.29 for Strong model (1.5B-FT) as expected.
Using fine-tuned Weak (0.5B-FT) and Strong (1.5B-FT) models, we generate weak labels from $\mathcal{D}_{WG}$ by comparing reward scores, as described in Sec. A.2, to train the Reward W2SG models. When the base 0.5B model is trained with weakly generated labels from the 0.5B model ‘Reward W2SG (0.5B-0.5B)’, it achieves a score of 233.01. This result surpasses its base model ‘Weak model (0.5B-Base)’ with a score of 211.52 but is worse than the directly fine-tuned version ‘Weak model (0.5B-FT)’ with a score of 228.29.
However, when the 1.5B model is trained with weakly generated labels from either the 0.5B model or the 1.5B model, the results are as follows: ‘Reward W2SG (0.5B-1.5B)’ achieves a score of 256.13, and ‘Reward W2SG (1.5B-1.5B)’ achieves 255.72. Both results surpass the fine-tuned 1.5B variant ‘Strong model (1.5B-FT)’ with a score of 228.29, as well as the fine-tuned 0.5B variant ‘Weak model (0.5B-FT)’ with a score of 246.52.
A.4 Debate Argument Analysis
We analyze the debate statements between two debaters with different capabilities. Tab. LABEL:tab:debate_comparison shows a comparison of statements between two separate debaters—GPT-4o as Debater 1 and GPT-4-Turbo as Debater 2, based on the story titled “Grandma Perkins and the Space Pirates”.
In the first round of the debate, the main difference is in their approaches to evidence. Debater 1 focuses on explicit textual evidence, citing specific quotes from the story. In contrast, Debater 2 relies heavily on inference. Debater 1 argues that the story describes only the return trip from Callisto to Earth, citing the lack of evidence for a round trip. Debater 2, however, emphasizes the word “back,” inferring that it implies a round trip.
In the second round, Debater 1 defends its position by challenging Debater 2’s reliance on implicit assumptions and the absence of explicit evidence. Debater 2 counters by asserting that the term “back” inherently implies a round trip but fails to provide additional textual evidence to substantiate this claim.
In the third round, Debater 1 highlights that Debater 2’s argument relies on assumptions rather than direct evidence. Debater 2 continues to argue that the implied meaning of “back” is sufficient to infer a round trip but does not address the lack of explicit confirmation in the story.
In short, Debater 1’s argument is more convincing to the judge because it is grounded in explicit textual evidence and direct quotes from the story. In contrast, Debater 2’s argument relies on assumptions about the term “back” without supporting textual evidence.
Additionally, we analyzed the debate when GPT-4o defended the incorrect answer and GPT-4-Turbo defended the correct one, and vice versa. The key difference lies in their approach: GPT-4o uses multi-step reasoning, incorporating multiple pieces of evidence and layers of analysis to support its argument, while GPT-4-Turbo focuses on a single, direct reason for its answer. Additionally, GPT-4o actively engages in defending its argument against alternative claims, whereas GPT-4-Turbo maintains a more straightforward assertion that its argument is correct.
Table 2: Full judge accuracy across various debater and judge setups in GPQA and QuALITY. All results are mean values from four independent runs. ‘Direct Access’ indicates whether the judge has direct access to the story. ‘# Correct’ represents the number of times the judge selects the correct answer, and ‘# Incorrect’ represents the number of times the judge selects the incorrect answer. ‘None’ is the baseline where the judge answers the question without debater. The results show that different capabilities between debaters influence the judge to identify the correct answer. For example, in GPQA, when the helpful debater is stronger (GPT-4o vs. GPT-4-Turbo), the judge’s accuracy improves to 55.42% with a GPT-4-Turbo judge and 65.75% with a GPT-4o judge. However, when the sneaky debater is stronger, performance drops to 52.00% with a GPT-4-Turbo judge and 51.92% with a GPT-4o judge. A similar trend is observed in QuALITY.
GPQA
[…]
GPT-4-Turbo GPT-4-Turbo GPT-4-Turbo 52.83% 158 142
GPT-4-Turbo GPT-4o GPT-4-Turbo 52.00% 156 144
GPT-4o GPT-4-Turbo GPT-4-Turbo 55.42% 166 134
GPT-4o GPT-4o GPT-4-Turbo 55.08% 165 135
None None GPT-4-Turbo 52.72% 158 142
GPT-4-Turbo GPT-4-Turbo GPT-4o 57.00% 171 129
GPT-4-Turbo GPT-4o GPT-4o 51.92% 156 144
GPT-4o GPT-4-Turbo GPT-4o 65.75% 197 103
GPT-4o GPT-4o GPT-4o 59.92% 180 120
None None GPT-4o 65.00% 195 105
QuALITY
[…]
GPT-4-Turbo GPT-4-Turbo GPT-4-Turbo 71.54% 67 27 X
GPT-4-Turbo GPT-4o GPT-4-Turbo 68.62% 65 29 X
GPT-4o GPT-4-Turbo GPT-4-Turbo 79.52% 75 19 X
GPT-4o GPT-4o GPT-4-Turbo 77.93% 73 21 X
None None GPT-4-Turbo 48.14% 45 49 X
None None GPT-4-Turbo 93.88% 88 6 O
GPT-4-Turbo GPT-4-Turbo GPT-4o 68.35% 64 30 X
GPT-4-Turbo GPT-4o GPT-4o 62.77% 59 35 X
GPT-4o GPT-4-Turbo GPT-4o 82.98% 78 16 X
GPT-4o GPT-4o GPT-4o 78.19% 74 20 X
None None GPT-4o 46.81% 44 50 X
None None GPT-4o 97.87% 92 2 O
Figure 5: Training accuracy on CosmosQA across varying model sizes (left:smallest to right:largest) and levels of noise in the training dataset (0.0 to 1.0). The larger model, Qwen2.5-7B-Instruct, shows a stronger tendency to overfit, learning even in noisy conditions, except in the range of 0.4 to 0.6, where labels are heavily randomized. In contrast, the smaller model, Qwen2.5-3B-Instruct, struggles to learn effectively when noise levels are between 0.3 and 0.7.
Table 5: CosmosQA evaluation accuracy when trained with different noise levels reveals that AI systems with large model parameters tend to overfit noisy data. As a result, their evaluation accuracy degrades rapidly (e.g., Qwen2.5) compared to smaller models (e.g., GPT-2).
Noise GPT-2
GPT-2
Medium
GPT-2
Large
Phi-3.5-Mini
Instruct
Qwen2.5-3B
Instruct
Qwen2.5-7B
Instruct
0.0 65.0% 69.5% 71.7% 83.8% 86.6% 87.2%
0.1 68.4% 69.4% 72.7% 81.3% 84.3% 86.0%
0.2 67.2% 66.6% 69.0% 81.8% 80.9% 81.1%
0.3 58.8% 62.0% 63.5% 77.8% 59.4% 76.3%
0.4 57.7% 57.3% 61.9% 76.8% 56.1% 56.4%
0.5 45.7% 46.9% 41.6% 47.2% 44.9% 45.6%
0.6 41.9% 42.7% 41.7% 25.2% 45.7% 43.4%
0.7 35.7% 40.4% 34.4% 23.4% 24.9% 23.6%
0.8 43.6% 34.1% 29.3% 21.0% 18.4% 18.2%
0.9 34.2% 30.8% 29.3% 17.9% 15.7% 15.2%
1.0 35.0% 30.5% 28.3% 16.2% 13.3% 12.1%
Table 8: Thinking advice prompt used to fill the [Thinking Advice] tag in debater prompts at each round. It consists of guidance on analyzing the math question, constructing logical arguments, critiquing the opponent’s reasoning, and refining responses over multiple debate rounds.
Table 9: The argument request prompt is inserted into the [Argument Request] tag in the debater prompts for each round. It is used to provoke the opposing debater into an argument.
Table 10: Debate judge prompt for GPQA debate. [ ] represents placeholders. Given a question, two possible answers (A and B), and a debate transcript, the judge’s task is to determine which answer is more convincing based on logical consistency, mathematical rigor, and clarity.
Table 11: Datasets used for LAIF evaluation. Both datasets are considered challenging enough to evaluate scalable oversight. For GPQA reasoning, we use the main set recommended by the authors for primary experiments. For QuALITY long reading comprehension, we filter the dataset using the same strategy as in (Khan et al., 2024) and divide it into two subsets: one for training and development, and the other for testing. The numbers in parentheses represent the number of unique stories in each subset, which are not shared across subsets.
Table 12: Template used to train the SFT model. [ ] represents placeholders. The training task is to learn how to generate explanations and predict the correct answer (True/False).
Table 13: Prompt used for explanation generation. [ ] represents the placeholder. Assistant generates [Generated Explanation] given the User message with story, question, and answer.
Table 14: Template used for W2SG experiment. [ ] represents placeholders. The training task is to learn to predict the correct answer (True/False).
Table 15: Overview of language models used in the paper, including their parameter sizes, instruction-tuning status, and whether they are proprietary. Models with an “O” in the Instruction Tuned column have been fine-tuned for following instructions, while those with an “X” in the Is Proprietary column are open-source.